OpenStudy (anonymous):

Find the length of the curve x=1+9t^2, y=3+6t^3, (0

OpenStudy (freckles):

hey

OpenStudy (anonymous):

can u help?

OpenStudy (freckles):

\[\text{ we have to use this } \int\limits_{a(t)}^{b(t)} \sqrt{(x')^2+(y')^2} dt \text{ or try to use } \\ \int\limits_{a(x)}^{b(x)} \sqrt{ 1+(y')^2} dx\]

OpenStudy (anonymous):

oh okay great

OpenStudy (freckles):

do you want to try to write y in terms of x?

OpenStudy (anonymous):

yes please

OpenStudy (freckles):

i'm asking that because if we try to write the x limits in terms of t limits we will end up with 4 numbers 2 of them being imaginary

OpenStudy (freckles):

this question is a tad weird for me it is 0<x<6 right?

OpenStudy (anonymous):

yea it says Find the length of the curve X=1+9t^2, and y+3+6t^3 (0 <t<6)

OpenStudy (anonymous):

that should clarify

OpenStudy (freckles):

oh that makes the problem tons easier you said 0<x<6

OpenStudy (anonymous):

yeah sorry

OpenStudy (freckles):

then just use that one parametric formula I gave you

OpenStudy (freckles):

\[\int\limits\limits_{0}^{6}\sqrt{((1+9t^2)')^2+((3+6t^3)')^2} dt\]

OpenStudy (anonymous):

i did but my answer wasn't right can u help me check the final answer again?

OpenStudy (freckles):

ok we can check you work what did you get for (1+9t^2)'? and for (3+6t^3)'?

OpenStudy (anonymous):

i got 461 for the final answer, but know thats wrong what did u get?

OpenStudy (freckles):

ok but what did you get for the derivatives of those two expressions I asked you above

OpenStudy (anonymous):

i did but just wanted to check my final answer first

OpenStudy (freckles):

is there anyway you can tell me what you got for the derivative of (1+9t^2) and (3+6t^3)

OpenStudy (freckles):

did you want me to check your problem?

OpenStudy (freckles):

In order for me to do that I need to see where you went wrong if anywhere

OpenStudy (anonymous):

yes i wanted to review the retriceer first then go over questions if possible

OpenStudy (anonymous):

18t and 18t^2 is the detractive

OpenStudy (anonymous):

i just like to see once i have the answer if i can work back and see my error it helps my learning

OpenStudy (freckles):

great! now inputting those in we get: \[\int\limits_{0}^{6}\sqrt{(18t)^2+(18t^2)^2} dt \\ \sqrt{18^2} \int\limits_0^6\sqrt{t^2+t^4} dt \\ 18 \int\limits_0^6 \sqrt{t^2+t^4} dt\] did you get to this part?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then do i solve from there?? because i got 29200?

OpenStudy (freckles):

notice both terms in the sqrt( ) have a t^2 in common \[\text{ since } t>0 \text{ then we can write }\\ 18 \int\limits_0^6 \sqrt{t^2} \sqrt{1+t^2} dt \\ \\ 18 \int\limits_0^6 t \sqrt{1+t^2} dt \\ \text{ now use substituion }\]

OpenStudy (freckles):

let u=1+t^2

OpenStudy (anonymous):

i have to go soon because my battery is about to die, could you share the answer?

OpenStudy (anonymous):

is mine right?

OpenStudy (anonymous):

please let me know mat 1%

OpenStudy (freckles):

I haven't done the problem. I was hoping we could do it together. If u=1+t^2 then du/dt=?

OpenStudy (freckles):

do you know power rule?

OpenStudy (freckles):

constant rule says (1)'=0 power rule says (t^2)=?

OpenStudy (freckles):

use wolframalpha if you just want answers